ramirez010202:

A store is having a sale on trail mix and jelly beans. For 9 pounds of trail mix and 7 pounds of jelly beans, the total cost is 53 . For 3 pounds of trail mix and 5 pounds of jelly beans, the total cost is 25. Find the cost for each pound of trail mix and each pound of jelly beans.

Supremesavage:

what do u think??

ramirez010202:

i honestly dont have a clue

Supremesavage:

ok so...

Supremesavage:

@sealake101

sealake101:

ok so what do u think would be the best step to do first?

ramirez010202:

i suck at math i really dont have even a bit of a guess

sealake101:

so we should probally find the total cost of the jelly beans

sealake101:

then the trail mix then u divide it to see how much each pound costs

sealake101:

ok?

jimthompson5910:

x = cost of 1 pound of trail mix y = cost of 1 pound of jelly beans We're told that "For 9 pounds of trail mix and 7 pounds of jelly beans, the total cost is 53" So we can form this equation 9x+7y = 53 Also, we're told that "For 3 pounds of trail mix and 5 pounds of jelly beans, the total cost is 25" leading to this equation 3x+5y = 25 The system of equations we need to solve is 9x+7y = 53 3x+5y = 25